Bonds issued by XYZ have a par value of $1000, were priced at $960 1 year ago and are priced at $940 today. The bonds pay annual coupons and just made a coupon payment. If the bonds had a percentage return over the past year ( from 1 year ago to today) of 17.00%, then what is the current yield of the bonds today?
Answer #1

Percentage return = (Ending value - begining value + coupons) / beginning value

0.17 = (940 - 960 + coupons) / 960

163.2 = 940 - 960 + coupons

183.2 =  coupons

Current yield = (Annual coupon / price) * 100

Current yield = (183.2 / 940) * 100

Current yield = 19.49%